### v2

# CHARTOM: A Visual Theory-of-Mind Benchmark for Multimodal Large Language Models

Abstract
We introduce CHARTOM, a visual theory-of-mind benchmark for multimodal large language models. CHARTOM consists of specially designed data visualizing charts. Given a chart, a language model needs to not only correctly comprehend the chart (the FACT question) but also judge if the chart will be misleading to a human reader (the MIND question). Both questions have significant societal benefits. We detail the construction of the CHARTOM benchmark including its calibration on human performance. We benchmarked leading LLMs as of late 2024—including GPT, Claude, Gemini, Qwen, Llama, and Llava—on the CHARTOM dataset and found that our benchmark was challenging to all of them, suggesting room for future large language models to improve.
1 Introduction
For AI to better assist humans, it must know not just factual truth but also how humans perceive the truth. The two could differ for many reasons such as information asymmetry or cognitive ability. If AI can detect the difference, it can help humans accordingly. The ability of an individual to reason about how others think (rather than the factual truth) is known as the theory of mind [34]. A classic example is the Sally-Anne test [43, 3]: Sally hides a marble in a basket and then leaves the room. Anne moves the marble to a box while Sally is away. Sally comes back. At this point, one can ask an observer two types of questions:
  • FACT question: Where is the marble?
  • MIND question: Where will Sally look for the marble?
There is a recent surge of interest in AI theory of mind, where large language models take the place of the observer. Several researchers found current AI competent at theory of mind tasks, performing near or surpassing human performance [22, 39], though some others remained skeptical [41, 37]. Concerned with limitations of AI theory of mind benchmarks, recent work also started to develop new benchmarks emphasizing conversations [20], causality [12], and actions [47]. However, these existing AI theory of mind benchmarks are all based on text comprehension.
Our theory of mind benchmark tests visual perception instead. Specifically, our focus is on misleading data visualizing charts, e.g. bar charts, line graphs, pie charts, etc. Charts are widely used for conveying quantitative information to human readers. They lend credibility to accompanying messages and hence have a strong impact on human decision-making. On the flip side, however, if charts are misleading, they can amplify the impact of misinformation [24, 33]. Figure 1 reproduces a misleading chart whose $y$-axis is inverted, which may give the impression that enacting Florida’s “stand your ground” law reduced gun deaths. Misleading charts are unfortunately prevalent in many areas of our lives. For example, a review of medical advertisements found that 1/3 of charts were misleading [8]. A review of charts in Korean news about the COVID-19 pandemic showed that about 30% of bar charts and 44% of pictorial charts were misleading [23]. A U.S.-based analysis of visuals that had been identified as misleading by human fact-checkers showed that more than half of these visuals were used as evidence for fake news about COVID-19 [5]. Because journalists often use charts to make narratives appear objective and to engage viewers emotionally [36], misleading charts can make an inaccurate story more deceptive [9, 42].
Figure 1: Reproduction of a misleading chart from [11], attributed to Reuters originally
As individuals increasingly rely on digital media to inform their opinions and decision-making processes, it is essential to ensure that the visual representations of data they encounter are accurate and trustworthy. Studying misleading charts can lead to AI with the ability to automatically detect them and alert readers, thus empowering people to make well-informed decisions based on reliable information, consequently promoting a more informed and discerning society.
1.1 Related Works
Existing computer vision research in charts have largely focused on chart comprehension. Classic approaches often employ an intermediate step of chart-to-data parsing, then use the data for downstream tasks such as visual question-answering, reasoning, and summarization. To this end, [6, 31] applied deep learning-based object detection methods to extract data, while others [1, 7, 29, 18, 38] applied semi-automated off-the-self OCR text recognition systems [4, 19]. Newer LLM-based approaches tend to be end-to-end, such as training the model on extensive pre-training datasets [28, 14, 30, 46], few-shot tuning methods which fine-tune pre-trained models with a small number of task-specific examples [10, 26, 21], and even zero-shot learning where the model makes predictions on tasks it has not been explicitly trained on [44, 17]. The newer approaches are performant, though occasionally still suffer from hallucination in data extraction and reasoning [16, 15].
However, we point out a significant shortcoming in chart comprehension: many misleading charts in fact contain the correct data. Take Figure 1 for example. Despite the unusual flipped $y$-axis, a meticulous reader (or AI) can indeed recover all the correct (year, deaths) data points from the chart. Therefore, ironically, a perfect chart comprehension system is not directly suited for judging how misleading a chart may be to humans. This is the same distinction between factual and theory of mind questions. We can ask:
  • A FACT question: What data is objectively presented in the chart?
  • A MIND question: What will a typical human reader perceive from the chart?
Current chart comprehension systems aim to solve the FACT question. To the best of our knowledge, the present paper is the first to address the MIND question in charts instead.
Our Contributions:
Towards this end, we make the following contributions:
  1. 1.
    We create a visual theory-of-mind benchmark CHARTOM (for CHARt Theory Of Mind) consisting of specially designed charts. The benchmark covers common variety of chart types and planted visual manipulations in those charts as shown in Table 1
  2. 2.
    We designed two types of questions for each chart: 1.) FACT question that are based on factual information present in charts, and, 2.) MIND question based on human theory of mind.
  3. 3.
    The key to the benchmark is to obtain MIND ground truth, i.e. “degree of misleadingness to humans”, for each chart that has been obtained via human experiments. This ground truth enables accessing whether an AI correctly answers the MIND questions.
  4. 4.
    We test the performance of leading LLMs in Claude, Gemini, GPT, Llama, LLava, and Qwen series as of Nov 2024 and report our findings. We observe that all LLMs perform poorly on both FACT and MIND questions indicating a big scope of improvement.
2 The CHARTOM Benchmark
Figure 2: A pair of non-manipulated(left) and manipulated(right) charts in CHARTOM with truncated y-axis manipulation.
In this section, we present our visual theory of mind benchmark CHARTOM. We first detail our design of charts to accommodate both FACT and MIND type of questions in the benchmark, then explain how we obtain the MIND ground truth with human experiments.
2.1 Design of the CHARTOM Benchmark
To explain our CHARTOM benchmark, let us start with an example.
An example. The pair of charts in Figure 2 share the same underlying data: (year=2018, dogs=250) and (year=2019, dogs=300). The left chart has a standard design which we call “non-manipulated”, while in the right chart we intentionally planted a potentially misleading manipulation: we truncated the $y$-axis range so it starts at 240 instead of 0. This is to create a perceptual effect: the 2019 bar now appears much taller than the 2018 bar. For each of the two charts, we ask the following FACT and MIND questions.
  • The FACT question is a chart comprehension question:
    The following graph shows the number of dogs adopted. The dogs adopted in 2018 eat 1 million bags of dog food in their lifetimes. How much do the dogs adopted in 2019 eat in their lifetimes?
  • The MIND question is a human performance question:
    Here is a chart we will present to typical university students and ask them the following question:
    [THE FACT QUESTION ABOVE]
    What fraction of typical university students do you predict will be misled by the chart when answering the question? First, give your prediction as a decimal number between 0 and 1, then justify your prediction in words.
The CHARTOM benchmark. Our entire CHARTOM benchmark follows the same design principle. The benchmark consists of 112 charts, covering five prevalent chart types: line, bar, pie, scatter plot, and map. The charts are manually designed to be visually tidy. We present selected charts from the benchmark in Table 1 and Appendix A.
The 112 charts come in 56 pairs (one pair is already shown in Figure 2). Each pair has a non-manipulated (suffix _1) and a manipulated (suffix _2) version, the latter incorporates one of many visualization fallacies suggested by the psychology literature, see the ‘manipulation’ column in Table 1. The manipulated version is intended to mislead casual human readers. Each pair of charts also comes with a FACT question and a MIND question. The FACT question is typically a simple word problem based on chart comprehension, see examples in Appendix A. There are three types of FACT questions: multiple-choice, free text entry, and sorting. Importantly, the true underlying data can still be recovered from either version if one is careful. This means that the FACT question has the same answer on both versions. The MIND question has the same format as shown earlier for all charts.
To summarize, the CHARTOM benchmark consists of
  • 112 charts: 56 pairs of non-manipulated and manipulated versions.
  • Each chart comes with a FACT question and the answer key.
  • Each chart comes with a MIND question, and the answer key is the Human Misleadingness Index (HMI, see Section 3.2) observed in human experiments.
To avoid web crawlers inadvertently adding the benchmark to future LLM training data, the CHARTOM benchmark is not online but is available upon request.
Non-Manipulated Manipulated
chart type group ID HMI manipulation type HMI
line G1 0.51 inconsistent x-axis 0.68
line G2 0.76 compressed y-axis 0.88
line G3 0.03 inverted y-axis 0.83
line G4 0.05 truncated y-axis 0.92
line G9 0.19 dual axes 0.84
bar G7 0.31 truncated y-axis 0.53
bar G8 0.76 compressed y-axis 0.69
bar G10 0.19 pictorial bars 0.27
bar G12 0.11 3D effect 0.48
scatter G5 0.02 logarithmic y-axis 0.63
scatter G6 0.02 swapped axis 0.73
map G15 0.02 inverted color scale 0.15
pie G13 0.25 3D effect 0.79
pie G14 0.30 3D effect + pop out 0.79
Table 1: MIND ground truth answers (HMI) score averaged across various manipulation groups in the CHARTOM benchmark. Each row is for non-manipulated and manipulated version of charts within a manipulation group. HMI stands for Human Misleadingness Index, a number between 0 and 1 where a larger value suggests the chart is more misleading to humans.
3 Obtaining Ground Truth Answers
3.1 Ground Truth for FACT Questions
The FACT questions are based on ground truth data. Their ground truth answers are easily computable. Recall both non-manipulated and manipulated versions by design have the same underlying data, and that data can be exactly recovered regardless of visual manipulation. Therefore, a perfect human or AI with proper reasoning can always obtain the same correct FACT answer on both chart versions. For example, in Figure 2 the correct FACT answer is $1.21\times\frac{300}{250}=$ 1.2 million bags.
Specifically, on FACT questions we suggest using accuracy, defined as the number of correct answers over the number of FACT questions. To define correctness, we consider the three types of FACT questions separately:
  1. 1.
    Multiple choice questions. An answer is correct if it precisely matches the answer key.
  2. 2.
    Ranking questions. A ranking answer is correct only if all items are ranked correctly.
  3. 3.
    Numerical (free text) questions. For numerical answers, the answer key provides the exact answer assuming perfect perception. But an exact match can be too strong an assumption for either humans or LLMs when they can enter any free-text number. To judge numerical answers, we suggest using a ±$\pm 10\%$ tolerance interval: if the answer key is $a$, then an LLM answer in the interval $[0.9a,1.1a]$ is deemed correct.
3.2 Ground Truth for MIND Questions
In contrast, the MIND question is trickier to judge. On the non-misleading chart in Figure 2(left) where the design is standard, one might still argue that the chart is not perfect and could be a little bit misleading to humans: the numbers are too small, there is no grid guidelines to help visual alignment, etc. Conversely, on the misleading chart (right) even if we intentionally planted the truncated $y$-axis because we presumed it to be misleading based on existing psychology literature (e.g. [27]), we do not know a priori the actual effect on human readers. To avoid these difficulties we propose the Human Misleadingness Index (HMI), a number between 0 and 1 for each chart which is the fraction of the human population that is misled by the chart. The correct answer to a MIND question is the HMI value.
To obtain HMI we conducted human experiments, which is expensive and time consuming but is the only proper method. Our human experiment involves 68 university students with IRB approval. We did not directly ask human participants the MIND question because we do not anticipate people to be able to give reliable answers—after all, many were themselves misled by some of our charts. Instead, we asked them the FACT question, which is designed so that users who are misled by the chart are likely to give wrong answers. We randomized the order of non-manipulated and manipulated charts in each pair when presenting them to human participants to minimize the order effect. We estimate the HMI of a chart by the percent of our study population who did not give a correct answer on the FACT question.
We argue that HMI is a surrogate measure for how misleading the chart is. This definition is motivated by two factors:
  1. 1.
    The arithmetic involved in solving the FACT questions is straightforward for the education level of our participants. Indeed on most non-manipulated charts, the HMI is small.
  2. 2.
    On many manipulated charts the HMI is higher compared to their non-manipulated chart counterpart. Since the only difference between the pair is the planted manipulation in data visualization, we can attribute the increase in HMI to the manipulation being more misleading to humans.
Figure 3: The HMI values of the 112 charts in the CHARTOM benchmark, sorted.
Figure 3 shows the HMI values in the CHARTOM benchmark. These values are sorted from small (not misleading to humans) to large (very misleading). In addition, each HMI value is color-coded to indicate whether the chart is the non-manipulated version (blue) or the manipulated version (red). Generally speaking, the non-manipulated charts have small HMI, and the manipulated ones have large HMI. But this depends on the type of manipulation. As the HMI values in Table 1 indicate, some manipulations greatly misled people: inconsistent x-axis, inverted y-axis, dual axes plots, 3D effect, etc. While interestingly, some other manipulations did not confuse people much: compressed y-axis and pictorial bars are benign. Such differences are of interest to cognitive science. Our human experiments are described in greater detail in [35].
4 Benchmarking Leading LLMs
We evaluated leading LLMs including Claude Sonnet 3.5 [2], Gemini 1.5 Pro [40], GPT-4o [32], LLaVA-OneVision [25], Llama-3.2-90B [13], Qwen2-VL-72B [45] on both FACT and MIND questions and present our findings below. When evaluating LLMs, we suggest the following criteria:
  • On FACT questions, we suggest reporting accuracy.
  • On the numerical part of the MIND questions, we suggest reporting mean absolute error: $\frac{1}{n}\sum_{i=1}^{n}|m_{i}-a_{i}|$ where $n$ is the number of MIND questions, $m_{i}$ is an LLM’s answer on the $i$th MIND question, and $a_{i}$ is the HMI answer key to that MIND question.
  • On the justification part of the MIND questions, we suggest qualitative comparisons to the “chart manipulations” file provided in the benchmark. That file records the single manipulation we planted to the manipulated (suffix _2) version of each chart pair. If a manipulated chart has a high HMI value, a capable LLM should identify the corresponding manipulation as an important cause of misleadingness.
The following are our main findings.
4.1 Like Humans, LLMs are Fooled by Manipulations
We asked LLMs to answer FACT questions on both non-manipulated and manipulated charts and record their response. Figure 4 shows the error rate of each LLM separately on non-manipulated charts (blue), manipulated charts (orange), and all charts (green). We make the following observations:
  1. 1.
    LLMs perform significantly worse on manipulated charts as compared to non-manipulated ones. This is evident from the difference in height of blue and orange bars for each LLM in Figure 4. We performed a $\mathcal{X}^{2}$ test that achieved a p-value $\leq 0.05$ for all LLMs. This suggests that, like humans, LLMs are fooled by manipulations.
  2. 2.
    Top LLMs show near or superhuman performance on FACT questions. We observe that top LLMs like Claude, Qwen, and GPT-4o have performed better or very close to humans (on average) while others like Llama, Gemini, and LLaVA have shown relatively poor performance.
  3. 3.
    Some LLMs fail to understand factual information in non-manipulated charts. While LLMs like Claude, Qwen, and GPT-4o excel on non-manipulated charts, they still make mistakes on 15-25% of cases. In contrast, models like Llama, Gemini, and LLaVA perform significantly worse indicating a big scope of improvement even on clean charts.
Figure 4: The error rate of humans and LLMs on FACT questions
4.2 Certain Manipulations are More Misleading Than Others
Figure 5: Log odds-ratio of making mistakes by humans and LLMs.
We compute the log odds ratio of making mistakes by humans and LLMs while answering FACT questions on non-manipulated vs. manipulation conditions, and show the results in Figure 5. The log odds-ratio is computed as follows:
$\log(OR)=\log\left(\frac{\hat{p}_{m}/(1-\hat{p}_{m})}{\hat{p}_{n}/(1-\hat{p}_{n})}\right)$ (1)
where $\hat{p}_{m},\hat{p}_{n}$ are the maximum likelihood estimates of making mistakes on manipulated and non-manipulated charts, respectively, by a population of agents. A higher odds-ratio suggests that manipulation is more misleading. We report the log odds ratio for each group and manipulation type.
To estimate $\hat{p}_{m}$ and $\hat{p}_{n}$, we fit separate mixed-effects logistic regression models for human and LLM responses. For LLMs, the model included fixed effects for group id and manipulation type, along with random intercepts for individual LLMs and question id within each group id. Similarly, for human responses, the model included fixed effects for group id and manipulation type, with random intercepts for human id and question id within each group. Estimated marginal means from these models were used to obtain $\hat{p}_{m}$ and $\hat{p}_{n}$ values which were then used for computing log-odds ratio.
We make the following observations from Figure 5:
  1. 1.
    Manipulations that effectively mislead humans tend to also deceive LLMs: There is a moderate to strong correlation (Kendall’s Tau score of 0.60) between the rankings of 14 groups based on their odds ratio in humans vs LLMs.
    • Top manipulation groups like line+manipulated x-axis, line+inverted y-axis, scatter + swapped axes, scatter+logarithmic y-axis are significantly misleading to both.
    • Bottom manipulations like bar+truncated y-axis, bar+compressed y-axis do not mislead humans. However, some of them like bar+pictorial bar and line+compressed y-axis, pie+3D still mislead LLMs, suggesting different modes of failure.
  2. 2.
    The two groups have significantly different scales of log odds ratio: The log odds ratio of LLMs on top manipulations in Figure 5 is on the order of $10^{20}$, compared to that of humans which is roughly on the order of $10^{5}$. This shows that when LLMs are misled, they are misled to a significant extent further. This again suggests different modes of failure.
4.3 LLMs Failed to Estimate Chart Misleadingness (HMI)
We asked each LLM the MIND question (what fraction of humans will be misled) on each chart. Table 2 reports the mean absolute error (MAE) between the LLM-predicted and ground truth HMI scores.
LLM Claude Gemini GPT-4o Llama LLaVA Qwen
MAE 0.40 $\pm$ 0.02 0.37 $\pm$ 0.02 0.35 $\pm$ 0.02 0.46 $\pm$ 0.02 0.32 $\pm$ 0.02 0.29 $\pm$ 0.02
Table 2: MAE of LLMs on MIND questions
The large MAE values suggest that the LLMs lacked certain visual theory-of-mind: they cannot reliably predict the extent to which humans are misled by various chart manipulations.
Figure 6: MIND Performance: LLM predicted HMI lacks the distinction shown in human ground truth with respect to high/low groups.
LLMs lack discriminative power for predicting HMI scores for high and low true HMI charts.
We further grouped all the charts in our benchmark dataset into high-HMI charts (charts with HMI0.5absent0.5$\geq$ 0.5≥ 0.5) and low-HMI charts (those with HMI<0.5). We computed the average predicted HMI by each LLM on both groups and plotted the results in Figure 6. For reference, we show the ground truth HMI on the left. We make the following observations:
  1. 1.
    LLMs show limited discriminative power on MIND questions: While all LLMs predict high-HMI charts to be more misleading than low-HMI ones on average, the difference is much smaller compared to the gap seen in human ground truth on the left of Figure 6. This indicates that LLMs lack a good discriminative power on MIND questions.
  2. 2.
    LLMs under/over estimate HMI of high and low HMI charts: As a corollary, we observe that LLMs consistently underestimate HMI for high misleading charts(orange bars of LLMs are consistently lower than orange ground truth HMI bar of human on the left) and overestimate the HMI for low misleading charts(blue bars follow the opposite trend instead).
5 Conclusion
We presented a new theory-of-mind benchmark called CHARTOM for LLMs based on visual chart comprehension. Our benchmark contained two types of questions: FACT questions based on factual information, and, MIND questions based on human theory of mind. We obtained the ground truth for MIND questions from our human experiments. Finally, we evaluated the performance of leading LLMs on both FACT and MIND questions and reported our findings. In summary, we observed that these LLMs perform poorly on both FACT and MIND questions indicating substantial room for improvement. We hope our benchmark will contribute to the development of future LLMs.
Acknowledgments. We thank Robert Hawkins for discussions on the theory of mind and Yea-Seul Kim and Yuhang Zhao for discussions on computer vision for chart comprehension. This work was supported by NSF IIS 2202457.
Appendix A: Additional Results
A FACT Results
We also analyze the performance of each LLM on different manipulation groups, report their performance in Figure 7 and make the following observations:
  1. 1.
    For manipulations like line + inverted y-axis and manipulated x-axis, all LLMs perform perfect on non-manipulated versions but are totally fooled on manipulated versions.
  2. 2.
    Some models like Llama, and Llava exhibit poor performance on even non-manipulated versions containing pie, line and bar charts highlighting that they lack the understanding of even standard charts and questions of these groups.
  3. 3.
    Surprisingly, most models perform poorly on both non-manipulated and non-manipulated versions of bar + compressed y-axis and pie chart with Claude performing the best among on prior and Qwen performing best on the later one.
Figure 7: Performance of LLMs on FACT questions for each manipulation group.
B MIND Results
We also show finer predictions made by LLMs on each manipulation group in Figure 8. Each black dot represents the ground truth HMI for the manipulated version(on the y-axis) and the non-manipulated version(on the x-axis). The corresponding HMI prediction made by LLM is shown with a colored dot joined by an arrow. We make the following observations:
  1. 1.
    For manipulation groups that are low misleading(black dots 2,8,10,15 in the bottom left of Figure 8), all LLMs consistently overestimate the misleadingness in both non-manipulated and manipulated version of the group. This is evident from the corresponding arrows getting stretched towards the top right.
  2. 2.
    For highly misleading manipulation groups (black dots 3,4,5,6 in the top left of Figure 8), all LLMs consistently underestimate the misleadingness in manipulated version and overestimate misleadingness in non-manipulated version. This is evident from the corresponding arrows getting stretched towards the bottom right.
Figure 8: Performance of different LLMs on MIND task.
C Logistic Regression Models for Estimating $\hat{p}$
To estimate the probabilities of making mistakes ($\hat{p}_{m}$, $\hat{p}_{n}$) for both LLMs and humans, we fit separate mixed-effects logistic regression models using the lme4 package in R.
LLM Responses.
The following model was used to estimate the probability of making error for large language models:
glmer(error ~ group_id:manipulation_type +
      (1 | llm) + (1 | group_id:question_id),
      data = data, family = "binomial")
Note that we have used group id and manipulation type as fixed effects and llm identity and question id of different examples within a particular group as mixed effect.
Human Responses.
A structurally similar model was used for human responses, replacing llm with student id:
glmer(error ~ group_id:manipulation_type +
      (1 | student_id) + (1 | group_id:question_id),
      data = human_data, family = "binomial")
Appendix B: Selected Charts in Table 1
[Uncaptioned image] [Uncaptioned image]
G4_Q1_1 G4_Q1_2
The following graph shows military spending. What is the trend of military spending from 2000 to 2012? 1. It has increased. 2. It has decreased. 3. It stayed the same.
[Uncaptioned image] [Uncaptioned image]
G3_Q1_1 G3_Q1_2
The following graph shows CD shipments. What is the trend of CD shipments from 2017 to 2020? 1. It has increased. 2. It has decreased. 3. It stayed the same.
[Uncaptioned image] [Uncaptioned image]
G9_Q1_1 G9_Q1_2
Two fictional countries Tomainia and Bacteria have been vying for military supremacy. The figure shows the number of tanks the countries had over the years. Which country had more tanks in 2021? 1. Tomainia 2. Bacteria
[Uncaptioned image] [Uncaptioned image]
G1_Q1_1 G1_Q1_2
As a stack of \$1 bills, a company’s corporate tax in 2014 was 1 yard high. How high is the company’s corporate tax in 2018 as a stack of \$1 bills?
[Uncaptioned image] [Uncaptioned image]
G2_Q1_1 G2_Q1_2
A company’s corporate tax in 2018 weighed 1 ton in gold. How many tons in gold does the company’s corporate tax weigh in 2022?
[Uncaptioned image] [Uncaptioned image]
G12_Q1_1 G12_Q1_2
The following graph shows the profit of company X in millions. Which year was company X most profitable? 1. 2018 2. 2022 3. 2018 and 2022 were the same.
[Uncaptioned image] [Uncaptioned image]
G7_Q1_1 G7_Q1_2
The following graph shows the number of dogs adopted. The dogs adopted in 2018 eat 1 million bags of dog food in their lifetimes. How much do the dogs adopted in 2019 eat in their lifetimes?
[Uncaptioned image] [Uncaptioned image]
G8_Q1_1 G8_Q1_2
The following graph shows the number of cats adopted. The cats adopted in 2017 need 1 ton of kitty litter pellets. How many tons of kitty litter pellets do the cats adopted in 2018 need?
[Uncaptioned image] [Uncaptioned image]
G10_Q1_1 G10_Q1_2
The graph shows how many bushels of apples were harvested this year for different kinds of apples. If 3 trucks are needed to transport all the Pink Lady apples, how many are needed to transport all the Honeycrisp apples?
[Uncaptioned image] [Uncaptioned image]
G6_Q1_1 G6_Q1_2
The following graph shows the relationship between car weight and miles per gallon. As car weight increases, what happens to the car’s miles per gallon? 1. Miles per gallon increases. 2. Miles per gallon decreases.
[Uncaptioned image] [Uncaptioned image]
G5_Q1_1 G5_Q1_2
The following graph shows the number of ebola cases in Sierra Leone. An election took place in September 2014. When did ebola cases increase more? 1. Before the election 2. After the election
[Uncaptioned image] [Uncaptioned image]
G15_Q1_1 G15_Q1_2
Which state has a higher foreign-born population, Wisconsin (WI) or Florida (FL)? 1. Wisconsin (WI) 2. Florida (FL)
[Uncaptioned image] [Uncaptioned image]
G13_Q1_1 G13_Q1_2
The figure below shows market share for different companies. Sort companies by market share from largest (1) to smallest (5). The largest company (1) — the smallest company (5)
[Uncaptioned image] [Uncaptioned image]
G14_Q1_1 G14_Q1_2
The figure below shows market share for different fast food restaurants. Based on this figure, sort restaurants by market share from largest (1) to smallest (5). The largest company (1) — the smallest company (5)